# Pushing LLMs to Their Logical Reasoning Bound: The Role of Data Reasoning Intensity

### v3

@@ Appendix C DRI Calculation Process and Re-Cognizing Optimization Algorithm
Appendix D Theoretical Foundations of DRI
D.1 Definitions of Core Concepts
D.1.1 Intrinsic Cognitive Cost $C(\mathcal{M})$
For a model $\mathcal{M}$, $C(\mathcal{M})$ represents the aggregate cost of its reasoning process, determined by a function of key model-intrinsic factors:
$C(\mathcal{M})=f(S,R,A)$
Here, $S$ denotes model scale, which encompasses parameters size and the depth of transformer layers, collectively reflecting the baseline resource demand of the model. $R$ refers to reasoning computational complexity, specifically the number of operations required for executing logical deduction steps such as multi-step inference and symbolic manipulation. $A$ stands for architectural constraints, including design features like attention mechanisms and expert selection strategies that influence the efficiency of the reasoning process.
D.1.2 Data Reasoning Potential $E(\mathcal{D})$
For a dataset $\mathcal{D}$, $E(\mathcal{D})$ quantifies the latent reasoning value embedded in its samples, defined as a function of critical data characteristics:
$E(\mathcal{D})=g(T,L,K)$
$T$ represents structured reasoning traces, which are the step-by-step logical chains present in sample annotations. $L$ denotes logical component density, measuring the number of atomic reasoning units (such as causal inference and conditional judgment) per unit length of reasoning text. $K$ indicates semantic coherence, reflecting the consistency and relevance between consecutive reasoning steps within the dataset.
D.2 Derivation of Equation (1)
The effective reasoning capability $\eta(\mathcal{M},\mathcal{D})$ is derived from the principle of resource-efficiency tradeoff. The core logic underlying this derivation is that $E(\mathcal{D})$ embodies the reasoning information that can be exploited by the model, meaning higher $E(\mathcal{D})$ tends to promote better reasoning performance. Conversely, $C(\mathcal{M})$ reflects the resource consumption required for the model to complete reasoning, so higher $C(\mathcal{M})$ may limit the effective utilization of data potential.
Based on this relationship, we hypothesize that $\eta$ is positively correlated with $E(\mathcal{D})$ and negatively correlated with $C(\mathcal{M})$, leading to the proportional form:
$\eta(\mathcal{M},\mathcal{D})\propto\frac{E(\mathcal{D})}{C(\mathcal{M})}$
The equality in Equation (1) is established by normalizing this proportionality to a dimensionless metric, where the specific scaling coefficient is context-dependent, varying with different model types and dataset domains.
D.3 Boundary Conditions
Equation (1) operates under the following implicit constraints:
  • Compatibility Range: The model $\mathcal{M}$ must have a minimum capacity to process the dataset $\mathcal{D}$, i.e., $C(\mathcal{M})\geq\kappa\cdot E(\mathcal{D})$, where $\kappa>0$ denotes the minimal model-to-data capacity ratio required for meaningful reasoning. For highly mismatched pairs (such as lightweight models processing ultra-complex reasoning data), $\eta$ loses interpretability.
  • Diminishing Returns: When $E(\mathcal{D})$ exceeds the reasoning boundary of $\mathcal{M}$—a threshold determined by the model’s maximum processing capacity—the growth of $\eta$ slows down due to inherent capacity limitations.
  • Interaction Effects: The separability of $E(\mathcal{D})$ and $C(\mathcal{M})$ is not absolute. Specialized models may exhibit higher $\eta$ for specific data structures, which is captured by context-dependent adjustments to the proportionality.
D.4 Theoretical Justification of DRI Components
To validate the rationality of the proposed DRI metrics, we analyze the relationship between each component in the quantification formulas and the actual reasoning difficulty.
For the logical intensity score $S_{\text{ctx}}$:
  • The term $|\mathcal{E}_{i}|\cdot\bar{D}^{2}$ reflects the structural complexity of reasoning. Logical expressions ($\mathcal{E}_{i}$) are the core carriers of reasoning logic, and their quantity directly affects the information processing load. The square of the average nesting depth ($\bar{D}^{2}$) is used because deeper nesting implies more nested logical operations (such as nested "if-then" structures), which exponentially increases the difficulty of parsing and deduction—consistent with the observation that complex structures in reasoning tasks create computational bottlenecks.
  • The counts of predicates ($|\mathcal{P}_{i}|$) and constants ($|\mathcal{C}_{i}|$) capture the richness of the reasoning elements. More unique predicates mean more types of relationships need to be processed, and more constants increase the burden of entity mapping, both of which are basic factors affecting reasoning difficulty.
For the reasoning intensity score $S_{\text{opt}}^{(l)}$:
  • The precondition term $|\mathcal{R}_{l}|\cdot\bar{D}_{l}^{2}$ quantifies the complexity of the initial logical assumptions. Similar to the structural term in $S_{\text{ctx}}$, it ensures that the difficulty of establishing preconditions (a key step in reasoning) is adequately reflected.
  • The step-wise term $\sum_{k=1}^{T_{l}}(1+\#\text{Operations}_{l,k})\cdot D_{l,k}^{2}$ considers both the number of reasoning steps ($T_{l}$) and the complexity of each step. The operator count ($\text{Operations}_{l,k}$) directly measures the logical operations (AND/OR/NOT) involved, and the square of nesting depth ($D_{l,k}^{2}$) again emphasizes the impact of structural complexity—aligning with the intuition that each step’s difficulty is determined by both its logical operations and structural depth.

### v4

@@ A.3 Future Work
Appendix B SLKD Calculation Process and Re-Cognizing Optimization Algorithm
Algorithm 1 illustrates the calculation process of our SLKD score, and Algorithm 2 demonstrates the implementation procedure of our Re-Cognizing Optimization strategy.
Appendix C Details for the Experiments
C.1 Experimental Settings
In this section, we conduct experiments on four logical reasoning benchmarks: Reclor (contextual reasoning) (Yu et al., 2020), LogicBench (deductive reasoning) (Parmar et al., 2024), LogiQA (workplace logic analysis) (Liu et al., 2020), and LogiQA2.0 (enhanced adversarial patterns) (Liu et al., 2023). For our experiments, we utilize two backbone models: LLaMA3.1-8B-Instruction (Grattafiori et al., 2024) and Qwen2.5-7B-Instruction (Qwen et al., 2025). We refer to them hereafter as LLaMA3.1-8B and Qwen2.5-7B. Additionally, we include GPT-4 (OpenAI, 2023) and DeepSeek-V3 (DeepSeek-AI, 2024) as reference models. For more implementation details, please refer to the Appendix C.
The released code for this work is available at https://github.com/HuzhouNLP/Logical_Structure_as_Knowledge.
C.2 Training Configuration
Compute Resources
C.3 Artifact License and Terms of Use.
Our work leverages existing public logical reasoning benchmarks (Reclor, LogiQA, LogiQA2.0, LogicBench) and open-source models (LLaMA3.1-8B, Qwen2.5-7B) for experiments, adhering to their original academic use permissions. The anonymized code for SLKD calculation and re-cognizing optimization is released via a public repository (https://anonymous.4open.science/r/Logical_Structure_as_Knowledge-81B1/) for non-commercial research purposes only. We do not claim ownership of any original dataset or model content, and all derived artifacts (e.g., SLKD-scored samples) are restricted to research use in line with the source materials’ terms.
C.4 Consistency with Intended Use.
The existing datasets employed in this work are originally designed for evaluating LLM logical reasoning capabilities, which aligns with our research goal of enhancing model reasoning via structured logical knowledge. Our re-cognizing optimization strategy uses these datasets to validate the effectiveness of SLKD, without repurposing them for unintended scenarios. For the artifacts we create (SLKD metric, optimization framework), we specify their intended use as research tools for data-centric LLM reasoning enhancement, ensuring compatibility with the original access conditions of source datasets.
C.5 Personally Identifying Information and Offensive Content.
The benchmark datasets used in our experiments are standard, curated NLP resources widely adopted for logical reasoning research, which do not contain intentional personally identifying information or offensive content. Our logical element extraction and SLKD calculation processes focus solely on decomposing textual logical structures (predicates, constants, logical expressions) and do not introduce, amplify, or expose sensitive attributes. All data processing steps prioritize preserving the original dataset’s safety and compliance, with no use of personally identifiable or offensive content in experiments.
C.6 Examples of Different Score Data
C.7 Reasoning Intensity Score Calculation Prompt
Prompt for Combinatorial Reasoning in MCQA Instructions: Please analyze the following multiple-choice question data. For each option, extract the relevant preconditions, define the deduction target, outline the deduction steps based on the provided Predicates, Constants, and Logical Expressions, and determine whether the option is correct according to the given answer index. Follow these specific rules: 1. The **Output** should be a JSON object containing only the ‘option_analysis‘ field, which is a list of analyses for each option. 2. For each option in ‘option_analysis‘, include the following fields: - **option_index**: The index of the option (starting from 0). - **option_text**: The text content of the option. - **preconditions**: A list of relevant preconditions from the Logical Expressions that pertain to the option. - **deduction_target**: The abstracted logical conclusion that the option is attempting to establish. - **deduction_steps**: A step-by-step logical deduction process from the preconditions to the deduction target. Each step should include: - **step**: The step number. - **task**: A description of what is being checked or inferred in this step. - **expression**: The logical expression used in this step, enclosed in backticks. Here, logical symbols like "implies" is represented as $\Rightarrow"$, "and" as $\land"$, "or" as $\lor"$, "not" as $\neg"$, "for all" as $\forall"$, "there exists" as $\exists"$ in LaTeX notation. - **result**: The outcome of this step, enclosed in backticks. - If a deduction step cannot proceed due to unsupported premises, indicate the failure and terminate further steps for that option. - **is_correct**: A boolean indicating whether the option is correct (true) or not (false). The values of ’Answer’ are ’0’, ’1’, ’2’, ’3’, where ’0’ represents the first option, ’1’ represents the second option, and so on. If the value of ’option_index’ is the same as the value of ’Answer’, then the value of ’is_correct’ is ’true’. 3. **Format Requirements**: - The output must strictly follow the JSON structure as shown in the example. - Ensure consistency in field naming and hierarchy. - Do not include any additional fields or information not specified in the example. 4. **Important Considerations**: - Only include preconditions that are directly relevant to the option being analyzed. - Maintain logical rigor in deduction steps, ensuring each step follows from the previous ones based on the preconditions. - Avoid including unrelated preconditions to minimize complexity and enhance clarity. Example: - **Context**: In rheumatoid arthritis, the body’s immune system misfunctions by attacking healthy cells in the joints causing the release of a hormone that in turn causes pain and swelling. This hormone is normally activated only in reaction to injury or infection. A new arthritis medication will contain a protein that inhibits the functioning of the hormone that causes pain and swelling in the joints. - **Question**: The statements above, if true, most strongly support which one of the following conclusions? - **Options**: 1. Unlike aspirin and other medications that reduce pain and swelling and that are currently available, the new medication would repair existing cell damage that had been caused by rheumatoid arthritis. 2. A patient treated with the new medication for rheumatoid arthritis could sustain a joint injury without becoming aware of it. 3. Joint diseases other than rheumatoid arthritis would not be affected by the new medication. 4. The benefits to rheumatoid arthritis sufferers of the new medication would outweigh the medication’s possible harmful side effects. - **Answer**: 1 - **Predicates**: - $\text{Attack}(x,y)$: Represents ’x’ attacking ’y’ (e.g., the immune system attacking healthy cells). - $\text{Release}(x,y)$: Represents ’x’ releasing ’y’ (e.g., the release of a hormone). - $\text{Cause}(x,y)$: Represents ’x’ causing ’y’ (e.g., the hormone causing pain and swelling). - $\text{Activate}(x,y)$: Represents ’x’ activating ’y’ (e.g., the hormone being activated by injury or infection). - $\text{Inhibit}(x,y)$: Represents ’x’ inhibiting ’y’ (e.g., the protein inhibiting the hormone). - $\text{Contain}(x,y)$: Represents ’x’ containing ’y’ (e.g., the medication containing a protein). - **Constants**: - ImmuneSystem: The body’s defense mechanism. - HealthyCells: Cells in the joints that are not diseased. - Hormone: A chemical messenger involved in causing pain and swelling. - Pain: A sensation caused by the hormone. - Swelling: A condition caused by the hormone. - Injury: A condition that normally activates the hormone. - Infection: A condition that normally activates the hormone. - ArthritisMedication: A new medication for treating arthritis. - Protein: A component of the medication that inhibits the hormone. - **Logical Expressions**: - $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ - $\text{Release}(\text{ImmuneSystem},\text{Hormone})$ - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - $\text{Activate}(\text{Injury},\text{Hormone})$ - $\text{Activate}(\text{Infection},\text{Hormone})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ Option Analysis 1. **Option Index**: 0 - **Option Text**: Unlike aspirin and other medications that reduce pain and swelling and that are currently available, the new medication would repair existing cell damage that had been caused by rheumatoid arthritis. - **Preconditions**: - $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ - $\text{Release}(\text{ImmuneSystem},\text{Hormone})$ - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - **Deduction Target**: $\text{Repair}(\text{ArthritisMedication},\allowbreak\text{HealthyCellsDamage})$ - **Deduction Steps**: - **Step 1**: - **Task**: Check if $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})$ implies $\text{Damage}(\text{HealthyCells})$. - **Expression**: $\text{Attack}(\text{ImmuneSystem},\text{HealthyCells})\Rightarrow\text{Damage}(\text{HealthyCells})$ - **Result**: Supported by context (immune system attacking healthy cells causes damage). - **Step 2**: - **Task**: Check if $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ and $\text{Inhibit}(\text{Protein},\text{Hormone})$ imply $\text{Repair}(\text{ArthritisMedication},\allowbreak\text{HealthyCellsDamage})$. - **Expression**: $\text{Contain}(\text{ArthritisMedication},\text{Protein})\land\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\text{Repair}(\text{ArthritisMedication},\allowbreak\text{HealthyCellsDamage})$ - **Result**: Not supported. The context only states that the protein inhibits the hormone, not that it repairs damage. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: $\text{Repair}(\text{ArthritisMedication},\allowbreak\text{HealthyCellsDamage})$ cannot be derived from the given preconditions. - **Is Correct**: False 2. **Option Index**: 1 - **Option Text**: A patient treated with the new medication for rheumatoid arthritis could sustain a joint injury without becoming aware of it. - **Preconditions**: - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Activate}(\text{Injury},\text{Hormone})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\exists\text{Patient},\text{Injury}:[\text{Sustain}(\text{Patient},\text{Injury})\land\text{Unaware}(\text{Patient},\text{Injury})]$ - **Deduction Steps**: - **Step 1**: - **Task**: Determine the effect of $\text{Inhibit}(\text{Protein},\text{Hormone})$ from the medication. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})$ - **Result**: Supported by context: The protein inhibits the hormone that causes pain and swelling. - **Step 2**: - **Task**: Analyze the implication of inhibiting the hormone on pain and swelling. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\neg\text{Cause}(\text{Hormone},\text{Pain})\land\neg\text{Cause}(\text{Hormone},\text{Swelling})$ - **Result**: Supported by context: If the hormone is inhibited, it cannot cause pain and swelling. - **Step 3**: - **Task**: Infer the patient’s awareness of injury when pain and swelling are absent. - **Expression**: $\neg\text{Cause}(\text{Hormone},\text{Pain})\land\neg\text{Cause}(\text{Hormone},\text{Swelling})\Rightarrow\text{Unaware}(\text{Patient},\text{Injury})$ - **Result**: Supported by context: Without pain and swelling, the patient may not be aware of sustaining an injury. - **Step 4**: - **Task**: Combine the above implications to conclude the deduction target. - **Expression**: $\exists\text{Patient},\text{Injury}:[\text{Sustain}(\text{Patient},\text{Injury})\land\text{Unaware}(\text{Patient},\text{Injury})]$ - **Result**: Deduction is valid based on the inhibited hormone preventing awareness of injury. - **Is Correct**: True 3. **Option Index**: 2 - **Option Text**: Joint diseases other than rheumatoid arthritis would not be affected by the new medication. - **Preconditions**: - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\forall x[\text{JointDisease}(x)\land x\neq\text{RheumatoidArthritis}\Rightarrow\neg\text{Affect}(\text{ArthritisMedication},x)]$ - **Deduction Steps**: - **Step 1**: - **Task**: Check if the context provides information about other joint diseases. - **Expression**: $\text{JointDisease}(x)\land x\neq\text{RheumatoidArthritis}$ - **Result**: Not supported. The context only discusses rheumatoid arthritis. - **Step 2**: - **Task**: Determine if there is any implication that the medication specifically targets rheumatoid arthritis. - **Expression**: $\text{Contain}(\text{ArthritisMedication},\text{Protein})\Rightarrow\text{SpecificEffect}(\text{RheumatoidArthritis})$ - **Result**: Not supported. The context does not specify that the protein exclusively affects rheumatoid arthritis. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: Cannot conclude that the medication does not affect other joint diseases. - **Is Correct**: False 4. **Option Index**: 3 - **Option Text**: The benefits to rheumatoid arthritis sufferers of the new medication would outweigh the medication’s possible harmful side effects. - **Preconditions**: - $\text{Cause}(\text{Hormone},\text{Pain})$ - $\text{Cause}(\text{Hormone},\text{Swelling})$ - $\text{Inhibit}(\text{Protein},\text{Hormone})$ - $\text{Contain}(\text{ArthritisMedication},\text{Protein})$ - **Deduction Target**: $\text{Benefit}(\text{ArthritisMedication})>\text{HarmfulSideEffect}(\text{ArthritisMedication})$ - **Deduction Steps**: - **Step 1**: - **Task**: Identify the benefits of the medication based on inhibiting the hormone. - **Expression**: $\text{Inhibit}(\text{Protein},\text{Hormone})\Rightarrow\text{Reduce}(\text{Pain})\land\text{Reduce}(\text{Swelling})$ - **Result**: Supported by context: The protein inhibits the hormone, which causes pain and swelling. - **Step 2**: - **Task**: Determine if the context provides information about harmful side effects. - **Expression**: $\text{HarmfulSideEffect}(\text{ArthritisMedication})$ - **Result**: Not supported. The context does not mention any side effects of the medication. - **Step 3**: - **Task**: Derivation fails. - **Expression**: Derivation cannot proceed. - **Result**: Cannot compare benefits and harmful side effects due to lack of information on side effects. - **Is Correct**: False Tips for Option Analysis - **Preconditions**: Only include logical expressions that are directly relevant to the option being analyzed. Avoid listing all possible preconditions. - **Deduction Steps**: Ensure each step logically follows from the previous one based on the preconditions. If a step cannot be completed due to insufficient support from the preconditions, indicate the failure and stop further deductions for that option. - **is_correct**: This field should be true only for the option that matches the answer index. All other options should be false. - **Format Consistency**: Maintain the same JSON structure and field naming conventions across all options to ensure uniformity and ease of data extraction. - **Logical Accuracy**: Ensure that all logical expressions and deductions accurately reflect the relationships defined by the predicates and constants. Your Task Analyze the following Input data and generate the option_analysis section as per the example above. Replace the xxx placeholders in the example with actual data derived from the input. Input Data input_data_here Please generate the option_analysis section based on the above input data.